OpenStudy (anonymous):

The proportion below shows values for y to x that form a direct variation. What is the constant of variation?

OpenStudy (anonymous):

6/2 = 24/8

OpenStudy (anonymous):

2 3 4 or 6

OpenStudy (anonymous):

2,3,4,or 6

OpenStudy (campbell_st):

do you have a calculate what is the answer to \[6 \div 2 =\] and \[24 \div 8 = \]

OpenStudy (anonymous):

3/3 1 whole

OpenStudy (campbell_st):

they are separate questions... 1. \[6 \div 2 = \] 2. \[24 \div 8 = \]

OpenStudy (anonymous):

3 and 3

OpenStudy (anonymous):

so 3 ?????

OpenStudy (campbell_st):

thats is... so what do you think the constant of variation is..?

OpenStudy (anonymous):

im new with Constant of Variation. :(

OpenStudy (campbell_st):

ok so this is what it looks like for direct variation y = kx k is the constant of variation so in your questions you have \[6 = 3 \times 2 ..... and ...... 24 = 3 \times 8 \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thnx

OpenStudy (campbell_st):

hope it helps